5. Calculate the drift speed of the conduction electrons in the copper wire of diameter 2.0 mm 

through which a steady current of 1.5 A is flowing. Density of copper = 8.96 X 103 Kg/m3, Molar mass of copper = 63.5 X 10-3 Kg/mol, NA = 6.023 X 1023 electrons / mol.


6. A spherical capacitor has inner radius 20.0 cm and outer radius 20.5 cm. What must be the plate area of a parallel plate capacitor with the same capacitance and plate separation?


7. The plates of a parallel plate capacitor each have an area of 0.40 m2 and are separated by a distance of 0.02 m. They are charged until the potential difference between the plates is 3000V. Find the capacitance, the charge on the plates and the potential energy stored.Also calculate the energy density



1
Expert's answer
2021-05-12T11:28:14-0400

5.


"I=envA\to v=\frac{I}{enA}=\frac{4IM}{e\pi d^2N_A\rho}="


"=\frac{4\cdot 1.5\cdot 63.5\cdot!0^{-3}}{1.6\cdot10^{-19}\cdot3.14\cdot 0.002^2\cdot6.023\cdot10^{23}\cdot8960}=3.5\cdot10^{-5}\ (m/s)" . Answer


6.


"4\pi\epsilon\epsilon_0\cdot \frac{r_1r_2}{r_2-r_1}=\frac{\epsilon\epsilon_0A}{r_2-r_1}\to"


"A=4\pi r_1r_2=4\cdot3.14\cdot 0.2\cdot0.205=0.51496\ (m^2)" . Answer


7.


"C=\frac{\epsilon_0A}{d}=\frac{8.85\cdot10^{-12}\cdot 0.4}{0.02}=177\cdot10^{-12}\ (F)" . Answer


"C=q/U\to q=CU=177\cdot10^{-12}\cdot 3000=531\cdot 10^{-9}\ (C)" . Answer


"W=CU^2/2=177\cdot10^{-12}\cdot3000^2/2=796.5\cdot10^{-6}\ (J)" . Answer


"\omega=W/V=796.5\cdot10^{-6}/(0.4\cdot0.02)=0.0996\ (J/m^3)" . Answer